OpenStudy (anonymous):

Find the equation of the tangent line to the curve at the point . The equation of this tangent line can be written in the form where

OpenStudy (tkhunny):

Have you considered the 1st Derivative?

OpenStudy (anonymous):

Find the equation of the tangent line to the curve y=4sin(x) at the point (Pi/6,2). The equation of this tangent line can be written in the form y=mx+b where m=? b=?

OpenStudy (anonymous):

yes the first derivative is y=4cox(x)

OpenStudy (amistre64):

m = y'(pi/6)

OpenStudy (amistre64):

and b = -mx1 + y1

OpenStudy (anonymous):

so would m just be 4cos(x)(pi/6)?

OpenStudy (amistre64):

m = 4cos(pi/6)

OpenStudy (anonymous):

ohhh okay

OpenStudy (amistre64):

if we take point slope form: y - y1 = m(x-x1) and solve for y ... y = mx - mx1 + y1 so y = mx + b is such that b=-mx1 + y1

OpenStudy (anonymous):

so b would be .1862006358

OpenStudy (amistre64):

maybe -2sqrt(3) pi/6 + 2 , the wolf likes it

OpenStudy (anonymous):

Thank you!

OpenStudy (amistre64):

youre welcome